OpenStudy (anonymous):

Using the formula P = 2(a + b), find a when P = 15 and b = 5. 2.5 5 8 10

OpenStudy (aravindg):

Just substitute P = 2(a + b) 15=2(a+5) solve for a

OpenStudy (sweetdreamtan):

a= 5/2 or 2.5